OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points (3,-1) and (-5,-1). Note: I thought this would be answered using the y=mx+b formula, but I am having trouble remembering if that is the correct way to figure this out or if it is another formula.

OpenStudy (phi):

slope is change in y divided by change in x

hero (hero):

It's y = mx + b but use it for each set of points to produce -1 = 3m + b -1 = -5m + b Then set isolate b b = -3m - 1 b = 5m - 1 Then set b = b 5m - 1 = -3m - 1 5m + 3m = 0 8m = 0 m = 0/8 m = 0 Therefore no slope But the fastest way to do it is to notice that the y coordinates are the same for both sets of points. Anytime that happens, it will result in a zero slope situation.

hero (hero):

When it comes to these kinds of problems, all you need is y = mx + b

hero (hero):

The only reason you need the other formulas is in case you're asked to use them on a test.

OpenStudy (phi):

I would say the slope is zero m=0 "no slope" sounds to close to "undefined slope"

hero (hero):

Sorry, oral typo. Zero slope instead of no slope. I was always taught that zero meant "no", but when it comes to slopes it gets "picky".

OpenStudy (anonymous):

It is picky and it is also confusing.